OpenStudy (anonymous):

Helen is helping plan a trip for three classes at her son's school. School rules say that classes on trips must have 1 adult for every 6 children. One has 26 children; and the third has 28 children. Each class has 1 teacher. How many OTHER adults are needed for the trip?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

thats not one of the choices

OpenStudy (anonymous):

28+26=54 54/6=9 9-2 teachers = 7

OpenStudy (anonymous):

28 students for one class + 26 students for the other= 54 students 54 students/6 per adult= 9 adults You already have to teachers so 54-12=42 42 student / 6 students per adult = 7 adults